OpenStudy (anonymous):

subtract. simplify removing a factor of 1 when possible 17fb/f^2-b^2 - f-b/f+b

jimthompson5910 (jim_thompson5910):

\[\Large \frac{17fb}{f^2-b^2} - \frac{f-b}{f+b}\] \[\Large \frac{17fb}{(f-b)(f+b)} - \frac{f-b}{f+b}\] \[\Large \frac{17fb}{(f-b)(f+b)} - \frac{(f-b)(f-b)}{(f-b)(f+b)}\] \[\Large \frac{17fb-(f-b)(f-b)}{(f-b)(f+b)}\] \[\Large \frac{17fb-(f^2-2fb+b^2)}{(f-b)(f+b)}\] \[\Large \frac{17fb-f^2+2fb-b^2}{(f-b)(f+b)}\] \[\Large \frac{-f^2+19fb-b^2}{(f-b)(f+b)}\] \[\Large \frac{-f^2+19fb-b^2}{f^2-b^2}\] So \[\Large \frac{17fb}{f^2-b^2} - \frac{f-b}{f+b}=\frac{-f^2+19fb-b^2}{f^2-b^2}\]

jimthompson5910 (jim_thompson5910):

if that answer doesn't work, then try \[\Large\frac{-f^2+19fb-b^2}{(f-b)(f+b)}\]